i need this with the p=2l+2w formula. Two rectangular solar panels are placed on the roof of a house. One panel is 4 ft wider than the other. If there combined width is 21 ft, what is the width of each panel
Answer by JoelSchwartz(130)

w1=width of the first rectangle
w2=width of the second rectangle
w1=4+w2
w1+w2=21
4+2w2=21
2w2=17
w2=8.5
w1=12.5